# Memorization: A Close Look at Books

### v1

@@ 1 Introduction
In this paper, we investigate the extraction of entire books from Llama 3 pretrained, and from Llama 3.1 models, both pretrained and instruction-tuned. For the instruction-tuned model, we employ Nasr et al.’s SFT-based technique. Books are important, because they are at the center of several copyright litigation cases. [^1] https://www.theguardian.com/books/2025/apr/04/us-authors-copyright-lawsuits-against-openai-and-
microsoft-combined-in-new-york-with-newspaper-actions
They are also technically interesting targets to extract, because they tend to be long and unique. An ideal extraction method would be able to auto-regressively extract entire books from an LLM trained on them given just their first $N$ tokens. While such extraction method does not [yet] exist, we were able to auto-regressively extract a version of “Alice’s Adventures in Wonderland” from Llama 3 pretrained that closely resembles the original. We were also able to obtain high reconstruction rates, although not auto-regressively, for many more books, with several Llama models. Moreover, we show how the popularity of books present in the training data, and therefore the likelihood of their duplication, affects their memorization by Llama. We conduct an analysis of memorization by examining the piece-wise reconstruction rates of full-length books sourced from Project Gutenberg, cross-referencing the results with the number of ratings in GoodReads. The following summarizes our experiments and findings:
@@ 2.1 Memorization in LLMs
Prior work has demonstrated that LLMs are capable of memorizing training data and susceptible to malicious extraction attacks Carlini et al. (2019, 2021); Thakkar et al. (2021); Ramaswamy et al. (2020); Lee et al. (2022); Zhang et al. (2023); Hayes et al. (2024). This memorization capability increases with the model size, the degree of duplication in the training data, and the length of the context prompt provided to the model Carlini et al. (2023); Kandpal et al. (2022). While many studies focus on open-source LLMs with accessible training datasets, some recent works have also proposed techniques to determine whether specific data have been used in training proprietary LLMs Chang et al. (2023); Ravichander et al. (2025). Nasr et al. proposed divergence attack and finetuning attack to extract training data from proprietary aligned models Nasr et al. (2025) Although these works provide valuable insights into data retention and memorization, they did not explore the reconstruction of entire works and the impact of data duplication, which is the focus of our work.
@@ 3.2 Datasets
We collect 32 English books from Project Gutenberg along two key dimensions: date added and popularity (see Table 1). The date of addition allows us to distinguish between books that were likely seen during training and those that were not. Project Gutenberg continues to grow through volunteer contributions, adding over 20 books in just the last 24 hours at the time of writing [^2] https://www.gutenberg.org/browse/recent/last1. Books added after Llama 3’s training cutoff are unlikely to have been included in the training data.
@@ 3.3 Data Extraction
Books #Chunks Min /Book Max /Book
@@ 3.4 Supervised Fine-Tuning
We fine-tune two variants of Llama 3.1 70B: pretrained abnd instruction-tuned. The samples were randomly chosen from 43 additional Gutenberg books not part of the extraction dataset. We run experiments with two distinct sample sizes: 500, and 1,000. Fine-tuning is performed on an NVIDIA RTX 6000 Ada GPU, leveraging the Unsloth  [^4] https://github.com/unslothai/unsloth, which facilitates efficient fine-tuning through quantization and Low-Rank Adaptation (LoRA). We use a learning rate of 2e-4, a batch size of 2, and train each model for one epoch.
@@ 3.5.2 Exp 2: Pretrained & Instruct STF
We use the same set of 9 books from the baseline experiment to investigate how different fine-tuning sample sizes affect the LLM’s memorization.
@@ 3.5.3 Exp 3: Expanded Study with STF-1000
We extend our extraction analysis to the instruct model fine-tuned on 1,000 samples. In this expanded experiment, we use all the books from the extraction dataset (Table 1).
@@ 3.6 Evaluation Metrics
We use a set of similarity metrics, including cosine similarity, Levenshtein distance, BLEU, Jaccard similarity, Sequence Matcher Similarity, and ROUGE-L.
@@ 4.1.1 Autoregressive Chunk Generation
The accompanying Figure  1, summarizes these results across all books and models for Jaccard similarity. Figures  7 and  8 displays results for BLEU and ROUGE-L similarity, respectively. Each bar represents the model’s entire generated output’s similarity to the whole ground truth text. Notably, the two post-2023 books are marked with an asterisk to highlight their addition after the Llama 3.1 knowledge cutoff date.
We find that Llama 3 70B exhibits the strongest memorization behavior. It achieves the highest similarity scores on nearly all books, particularly on widely duplicated texts such as Alice in Wonderland and The Time Machine, supporting the hypothesis that more popular books—likely duplicated across training corpora—are more easily regurgitated by less aligned models.
@@ 4.2 Exp 2: Supervised Fine-Tuning
As shown in the figure, books with higher number of ratings generally achieve significantly higher median Jaccard similarity scores compared to books with lower number of ratings. The correlation coefficient is 0.5, which is indicative of a fairly strong positive correlation. This correlation suggests that higher popularity may be associated with greater availability and duplication on the internet. The three books with the highest reconstruction rates are The Communist Manifesto (0.95), Alice’s Adventures in Wonderland (0.91), and Romeo and Juliet (0.76).
@@ 5 Analysis of Weight Updates
Figure 5 shows a histogram built from the concatenated set of values of all 560 $W_{\mathrm{rel}}$ matrices. The distribution clearly reveals that the vast majority of updates are relatively small in magnitude when compared to the original weights. Only about $\sim 14\%$ of the original weights are receiving a boost greater than only $1\%$, and a mere $\sim 0.15\%$ are updated by more than 100%. These results suggest that only sparse and highly localized updates are sufficient to make the instruct network start remembering documents that were used in its training set.
@@ 6 Conclusion
Appendix A Appendix
@@
A.1 Additional Results
A.1.1 Autoregressive Generation Experiments
This section presents the full set of similarity metrics used to evaluate autoregressive generation, including Jaccard, ROUGE-L, and BLEU.
@@ A.1.1 Autoregressive Generation Experiments
Figure 10: Median BLEU scores for books of varying popularity levels extracted from Llama 3.1 instruct SFT 1000 samples. Books from the pre-training collection (pre) and post-training collection (post) are indicated by blue and red markers, respectively.
Figure 11: Median Cosine Similarity scores for books of varying popularity levels extracted from Llama 3.1 instruct SFT 1000 samples. Books from the pre-training collection (pre) and post-training collection (post) are indicated by blue and red markers, respectively.
Figure 12: Median Levenshtein scores for books of varying popularity levels extracted from Llama 3.1 instruct SFT 1000 samples. Books from the pre-training collection (pre) and post-training collection (post) are indicated by blue and red markers, respectively.
Figure 13: Median ROUGE-L scores for books of varying popularity levels extracted from Llama 3.1 instruct SFT 1000 samples. Books from the pre-training collection (pre) and post-training collection (post) are indicated by blue and red markers, respectively.
Figure 14: Median ROUGE-L scores for books of varying popularity levels extracted from Llama 3.1 instruct SFT 1000 samples. Books from the pre-training collection (pre) and post-training collection (post) are indicated by blue and red markers, respectively.
Appendix B Analysis of Weight Updates

### v2

@@ 1 Introduction
In this paper, we investigate the extraction of entire books from Llama 3 pretrained, and from Llama 3.1 models, both pretrained and instruction-tuned. For the instruction-tuned model, we employ Nasr et al.’s SFT-based technique. Books are important, because they are at the center of several copyright litigation cases. [^1] https://www.nytimes.com/2023/12/27/business/media/new-york-times-open-ai-microsoft-lawsuit.html?smid=url-share They are also technically interesting targets to extract, because they tend to be long and unique. An ideal extraction method would be able to auto-regressively extract entire books from an LLM trained on them given just their first $N$ tokens. While such extraction method does not [yet] exist, we were able to auto-regressively extract a version of “Alice’s Adventures in Wonderland” from Llama 3 pretrained that closely resembles the original. We were also able to obtain high reconstruction rates, although not auto-regressively, for many more books, with several Llama models. Moreover, we show how the popularity of books present in the training data, and therefore the likelihood of their duplication, affects their memorization by Llama. We conduct an analysis of memorization by examining the piece-wise reconstruction rates of full-length books sourced from Project Gutenberg, cross-referencing the results with the number of ratings in GoodReads. The following summarizes our experiments and findings:
@@ 2.1 Memorization in LLMs
Prior work has demonstrated that LLMs are capable of memorizing training data and susceptible to malicious extraction attacks Carlini et al. (2019, 2021); Thakkar et al. (2021); Ramaswamy et al. (2020); Lee et al. (2022); Zhang et al. (2023); Hayes et al. (2024). This memorization capability increases with the model size, the degree of duplication in the training data, and the length of the context prompt provided to the model Carlini et al. (2023); Kandpal et al. (2022). While many studies focus on open-source LLMs with accessible training datasets, some recent works have also proposed techniques to determine whether specific data have been used in training proprietary LLMs Chang et al. (2023); Ravichander et al. (2025). Nasr et al. proposed divergence attack and finetuning attack to extract training data from proprietary aligned models Nasr et al. (2025). Zhao et al. use partial information probing, providing LLMs with excerpts from copyrighted texts and prompting them to complete the passages, in order to assess the extent to which LLMs can reproduce copyright-protected content Zhao et al. (2024). Although these works provide valuable insights into data retention and memorization, they did not explore the reconstruction of entire works and the impact of data duplication, which is the focus of our work.
While both Karamolegkou1 et al. Karamolegkou et al. (2023) and our work investigate the relationship between content popularity and memorization in LLMs, the prior work primarily quantifies verbatim memorization using the length of the longest common subsequence between generated and reference texts. In contrast, our study focuses on the feasibility of reconstructing entire books from LLMs, systematically evaluating extraction rates across both popular and obscure works to understand the boundaries of model memorization.
@@ 3.2 Datasets
We collect 32 English books (Table 5) from Project Gutenberg along two key dimensions: date added and popularity (see Table 1). The date of addition allows us to distinguish between books that were likely seen during training and those that were not. Project Gutenberg continues to grow through volunteer contributions, adding over 20 books in just the last 24 hours at the time of writing [^2] https://www.gutenberg.org/browse/recent/last1. Books added after Llama 3’s training cutoff are unlikely to have been included in the training data.
@@ 3.3 Data Extraction
Books #Chunks
Min
/Book
Max
/Book
@@ 3.4 Supervised Fine-Tuning
We fine-tune two variants of Llama 3.1 70B: pretrained and instruction-tuned. The samples are randomly chosen from 43 additional Gutenberg books (Table 4) not part of the extraction dataset. We run experiments with two distinct sample sizes: 500, and 1,000. Fine-tuning is performed on an NVIDIA RTX 6000 Ada GPU, leveraging the Unsloth [^4] https://github.com/unslothai/unsloth, which facilitates efficient fine-tuning through quantization and LoRA. We use a learning rate of 2e-4, a batch size of 2, and train each model for one epoch.
@@ 3.5.2 Exp 2: Pretrained & Instruct STF
We use the same set of 9 books from the baseline experiment to investigate how different fine-tuning sample sizes [500, 1000] affect the LLM’s memorization.
@@ 3.5.3 Exp 3: Expanded Study with STF-1000
We extend our extraction analysis to the instruct model fine-tuned on 1,000 samples. In this expanded experiment, we use all 32 books from the extraction dataset (Table 5).
@@ 3.6 Evaluation Metrics
We use a set of similarity metrics, including cosine similarity, Levenshtein distance, BLEU, Jaccard similarity, Sequence Matcher Similarity, and ROUGE-L to measure the extraction results.
@@ 4.1.1 Autoregressive Chunk Generation
The accompanying Figure 2, summarizes these results across all books and models for Jaccard similarity. Figures 9 and 8 displays results for BLEU and ROUGE-L similarity, respectively. Each bar represents the model’s entire generated output’s similarity to the whole ground truth text. Notably, the two post-2023 books are marked with an asterisk to highlight their addition after the Llama 3.1 knowledge cutoff date.
We find that Llama 3 70B exhibits the strongest memorization behavior. It achieves the highest similarity scores on widely popular texts such as Alice in Wonderland and The Time Machine, supporting the hypothesis that more popular books—likely duplicated across training corpora—are more easily regurgitated by less aligned models.
@@ 4.2 Exp 2: Supervised Fine-Tuning
As shown in the figure, books with higher number of ratings generally achieve significantly higher median Jaccard similarity scores compared to books with lower number of ratings. The correlation coefficient is 0.5, which is indicative of a fairly strong positive correlation. This correlation suggests that higher popularity may be associated with greater availability and duplication on the internet. The three books with the highest reconstruction rates are The Communist Manifesto (0.95), Alice’s Adventures in Wonderland (0.91), and Romeo and Juliet (0.76). See Appendix B for more details.
@@ 5 Analysis of Weight Updates
[…]
@@ 6 Conclusion
7 Limitations
Our study provides initial evidence on the extent of book extraction from LLMs, but several limitations should be noted, which in turn suggest directions for future research.
First, this study is limited to the Llama 3.x family of models, specifically the 70B parameter variants. While this focus enables a detailed examination of memorization and extraction within a widely used model, it restricts the generalizability of our findings to other model families and architectures. We did not investigate scaling effects or compare across different model sizes, as prior work  Carlini et al. (2023) has consistently demonstrated that memorization capacity increases with model size.
Second, our evaluation primarily targets books that are publicly available, particularly those released on Project Gutenberg prior to the Llama 3.x knowledge cutoff date. Future work could extend this analysis to books released after the cutoff date, as well as to widely known but copyright-protected works that are not part of the Project Gutenberg collection, such as A Farewell to Arms and the Harry Potter series.
Finally, we use book popularity, as measured by Goodreads ratings, as a proxy for the likelihood of duplication in the training data. While this approach is practical for published books, it may not generalize to other types of content, such as news articles or academic papers. For these domains, alternative metrics (e.g., number of downloads or citations) may be required, but their suitability as proxies for training data frequency remains to be validated.
Appendix A Book Details
The following tables provide detailed information for all books used in this study, including their titles, number of Goodreads ratings, authors, and initial Project Gutenberg release dates. The books are sorted in descending order of popularity. Table 4 lists the 43 books used for fine-tuning, while Table 5 lists the 32 books used for testing.
Ratings Book Names Author First Added On
5630463 The Great Gatsby F. Scott Fitzgerald January 17, 2021
4539830 Pride and Prejudice Jane Austen June 1, 1998
2199001 Jane Eyre: An Autobiography Charlotte Brontë March 1, 1998
1929915 Wuthering Heights Emily Brontë December 1, 1996
1307593 Adventures of Huckleberry Finn Mark Twain June 29, 2004
1293875 Metamorphosis Franz Kafka August 17, 2005
1250187 Sense and Sensibility Jane Austen September 1, 1994
1141312 The Odyssey Homer April 1, 1999
999417 Crime and Punishment Fyodor Dostoyevsky March 28, 2006
988702 A Tale of Two Cities Charles Dickens January 1, 1994
984922 The Adventures of Tom Sawyer, Complete Mark Twain July 1, 2004
975420 The Count of Monte Cristo Alexandre Dumas, Auguste Maquet January 1, 1998
886396 A Christmas Carol in Prose; Being a Ghost Story of Christmas Charles Dickens August 11, 2004
848348 Great Expectations Charles Dickens July 1, 1998
742250 Persuasion Jane Austen February 1, 1994
631406 The Strange Case of Dr. Jekyll and Mr. Hyde Robert Louis Stevenson June 27, 2008
539036 Heart of Darkness Joseph Conrad January 9, 2006
485278 The Iliad Homer July 1, 2004
392898 The Importance of Being Earnest: A Trivial Comedy for Serious People Oscar Wilde March 1, 1997
364727 The Prince Niccolò Machiavelli February 11, 2006
361107 The Brothers Karamazov Fyodor Dostoyevsky February 12, 2009
352180 War and Peace graf Leo Tolstoy April 1, 2001
328677 An Anglo-Saxon Epic Poem J. Lesslie Hall (translator) July 19, 2005
328025 The Yellow Wallpaper Charlotte Perkins Gilman November 1, 1999
317240 The Adventures of Sherlock Holmes Arthur Conan Doyle March 1, 1999
217766 Grimms’ Fairy Tales Jacob Grimm, Wilhelm Grimm April 1, 2001
217052 The Republic Plato October 1, 1998
166045 Thus Spake Zarathustra: A Book for All and None Friedrich Wilhelm Nietzsche December 1, 1999
136898 Ulysses James Joyce July 1, 2003
130667 Narrative of the Life of Frederick Douglass, an American Slave Frederick Douglass January 12, 2006
106656 Beyond Good and Evil Friedrich Wilhelm Nietzsche August 1, 2003
69676 The Confessions of St. Augustine Bishop of Hippo Saint Augustine June 1, 2002
50681 Leviathan Thomas Hobbes May 1, 2002
48030 A Modest Proposal Jonathan Swift October 1, 1997
46229 Cranford Elizabeth Cleghorn Gaskell January 1, 1996
43921 The Souls of Black Folk W. E. B. Du Bois January 1, 1996
38189 Walden, and On The Duty Of Civil Disobedience Henry David Thoreau January 1, 1995
23339 Second Treatise of Government John Locke January 1, 2005
21501 The King in Yellow Robert W. Chambers July 1, 2005
2726 The Letters of Jane Austen Jane Austen February 12, 2013
2548 The Works of Edgar Allan Poe — Volume 2 Edgar Allan Poe April 1, 2000
1277 The Adventures of Roderick Random T. Smollett May 1, 2003
65 The Devil is an Ass Ben Jonson October 7, 2015
Table 4: The fine-tuning dataset, consisting of 43 books from Project Gutenberg released before December 2023, sorted by the number of Goodreads ratings.
Ratings Book Names Author First Added On
2,735,023 Romeo and Juliet William Shakespeare November 1, 1998
831,152 Siddhartha Hermann Hesse February 1, 2001
546,286 The Time Machine H. G. Wells October 2, 2004
492,129 The Wonderful Wizard of Oz L. Frank Baum October 12, 2013
413,400 Alice’s Adventures in Wonderland Lewis Carroll June 27, 2008
362,694 Peter Pan J. M. Barrie June 25, 2008
290,524 Candide Voltaire November 27, 2006
219,332 The Tempest William Shakespeare October 26, 2007
198,834 Notes from the Underground Fyodor Dostoyevsky July 1, 1996
183,818 The Communist Manifesto Karl Marx, Friedrich Engels January 25, 2005
169,009 The Sign of the Four Arthur Conan Doyle March 1, 2000
147,072 The Legend of Sleepy Hollow Washington Irving June 27, 2008
140,217 Through the Looking-Glass Lewis Carroll June 25, 2008
129,512 The Island of Doctor Moreau H. G. Wells October 14, 2004
48,922 Just So Stories Rudyard Kipling August 1, 2001
19,734 Ethics Benedictus de Spinoza February 1, 2003
7,582 The Aesop for Children Aesop December 2, 2006
6,071 The Secret of the Caves* Franklin W. Dixon February 7, 2025
787 Simple Sabotage Field Manual United States. Office of Strategic Services August 4, 2008
344 The First Book of Adam and Eve Rutherford Hayes Platt January 1, 1996
138 The Philippines a Century Hence Austin Craig April 18, 2011
22 The Emma Gees Herbert W. McBride February 24, 2007
16 Bab Ballads and Savoy Songs W. S. Gilbert March 15, 2005
5 Dragon Moon* Henry Kuttner January 28, 2025
4 The Hallowell Partnership Katharine Holland Brown October 14, 2012
2 Judas Ram Sam Merwin January 27, 2016
2 Rosin the Beau Laura Elizabeth Howe Richards December 24, 2008
0 Christina and the Boys* Amy Le Feuvre February 10, 2025
0 Pegasus* J. F. C. Fuller January 30, 2025
0 A girl and her ways* Amy Le Feuvre January 28, 2025
0 Upside Down or Backwards W. C. Tuttle December 20, 2021
0 Science and Medieval Thought T. Clifford Allbutt February 21, 2012
Table 5: The testing dataset, consisting of 32 books released both before and after the cutoff date, sorted by the number of Goodreads ratings. * denotes a book released after the knowledge cutoff date (December 2023).
Appendix B Additional Results
@@
B.1 Autoregressive Generation Experiments
This section presents the full set of similarity metrics used to evaluate autoregressive generation, including Jaccard ( Figure 7), ROUGE-L (Figure 8), and BLEU (Figure 9).
@@ B.1 Autoregressive Generation Experiments
B.2 Llama 3.1-instruct SFT-1000
Median similarity scores for books extracted from Llama 3.1-instruct SFT-1000 samples are presented. Books from the pre-training collection (pre) and post-training collection (post) are indicated by blue and red markers, respectively. The exact median similarity scores are listed in Table 6. Visualizations of these results are shown in Figures 1014.
Ratings Book Names Median Jaccard Median Cosine Median Levenshtein Median Sequence Matcher Median BLEU Median ROUGE-L
2,735,023 Romeo and Juliet 0.762 0.906 0.931 0.957 0.708 0.941
831,152 Siddhartha 0.233 0.342 0.413 0.494 0.175 0.375
546,286 The Time Machine 0.250 0.345 0.432 0.506 0.228 0.392
492,129 The Wonderful Wizard of Oz 0.333 0.456 0.496 0.593 0.316 0.500
413,400 Alice’s Adventures in Wonderland 0.913 0.962 0.944 0.963 0.931 0.978
362,694 Peter Pan 0.182 0.257 0.356 0.427 0.108 0.292
290,524 Candide 0.267 0.362 0.433 0.536 0.192 0.419
219,332 The Tempest 0.370 0.467 0.535 0.638 0.348 0.587
198,834 Notes from the Underground 0.235 0.301 0.402 0.478 0.168 0.364
183,818 The Communist Manifesto 0.952 0.980 0.961 0.971 0.953 0.980
169,009 The Sign of the Four 0.237 0.311 0.420 0.496 0.191 0.367
147,072 The Legend of Sleepy Hollow 0.407 0.526 0.597 0.667 0.435 0.585
140,217 Through the Looking-Glass 0.458 0.566 0.651 0.724 0.485 0.652
129,512 The Island of Doctor Moreau 0.158 0.236 0.321 0.387 0.046 0.263
48,922 Just So Stories 0.188 0.292 0.376 0.453 0.113 0.333
19,734 Ethics 0.343 0.456 0.488 0.591 0.279 0.500
7,582 The Aesop for Children 0.179 0.288 0.336 0.424 0.053 0.298
6,071 The Secret of the Caves* 0.103 0.186 0.263 0.333 0.014 0.196
787 Simple Sabotage Field Manual 0.114 0.156 0.268 0.334 0.018 0.195
344 The First Book of Adam and Eve 0.229 0.343 0.383 0.468 0.120 0.346
138 The Philippines a Century Hence 0.111 0.217 0.262 0.313 0.012 0.186
22 The Emma Gees 0.108 0.187 0.262 0.314 0.013 0.182
16 Bab Ballads and Savoy Songs 0.100 0.167 0.304 0.387 0.015 0.200
5 Dragon Moon* 0.088 0.160 0.244 0.309 0.012 0.170
4 The Hallowell Partnership 0.079 0.130 0.244 0.304 0.011 0.160
2 Judas Ram 0.088 0.145 0.246 0.304 0.012 0.163
2 Rosin the Beau 0.108 0.157 0.261 0.317 0.012 0.174
0 Christina and the Boys* 0.093 0.151 0.252 0.321 0.012 0.174
0 Pegasus* 0.111 0.204 0.261 0.329 0.014 0.196
0 A girl and her ways* 0.098 0.141 0.252 0.310 0.012 0.174
0 Upside Down or Backwards 0.073 0.137 0.242 0.306 0.011 0.158
0 Science and Medieval Thought 0.100 0.216 0.257 0.309 0.013 0.186
Table 6: Median similarity scores for books extracted from Llama 3.1-instruct SFT-1000 samples. * denotes a book released after the knowledge cutoff date (December 2023).
Figure 10: Median BLEU scores
Figure 11: Median Cosine Similarity scores
Figure 12: Median Levenshtein scores
Figure 13: Median ROUGE-L scores
Figure 14: Median Sequence Matcher scores
Appendix C Analysis of Weight Updates